# Anisotropic Spin Ice on a Breathing Pyrochlore Lattice

### v1

@@ Abstract
Spin ice systems represent a prime example of constrained spin systems and exhibit rich low-energy physics. In this study, we explore how introducing a tunable anisotropic spin coupling to the conventional Ising spin ice Hamiltonian on the breathing pyrochlore lattice affects the ground state properties of the system. Significant changes are observed in the ground state structure, reflected in the spin structure factor and in a reduction of residual entropy at low temperatures. We theoretically uncover a rich phase diagram by varying the anisotropy and demonstrate how this modification reduces the ground state degeneracy across different phases. Numerical simulations reveal that, at sufficiently low temperatures, the system either undergoes a crossover into a constrained spin ice manifold, characterized by an entropy density that drops below the Pauling entropy of conventional spin ice, or a phase transition into a symmetry-broken state, depending on the perturbations. Additionally, we compute the spin structure factors for the anisotropic model and compare these results to analytical predictions from a large-$N$ expansion, finding good agreement. This work develops the understanding of spin ice in anisotropic limits, which may be experimentally realized by strain, providing, among others, key signatures in entropy and specific heat.
Anisotropic Breathing Ising Spin Ice
@@ I Introduction
While the simple nearest-neighbor spin-ice model is by now well understood [2, 6, 7, 8], several questions arise when comparing it to experiments. Experimental studies of Dy2Ti2O7 have revealed, for instance, that thermally equilibrated samples, cooled over long timescales, exhibit an entropy drop below the celebrated Pauling entropy, a phenomenon that remains under active investigation [9, 10]. One possible explanation is the formation of a quantum spin liquid, where the true ground state would be a coherent superposition of the classical configurations, thereby lifting the extensive degeneracy [11, 12, 13]. Another explanation for the release of residual entropy is based on further neighbor or long-range interactions present in spin ice materials. These naturally change the degeneracy of the ground state and lead to an ordered state at low energies [14, 15, 16]. An alternative approach to release the residual entropy has been shown to be through strain engineering. Through theoretical [17, 18] and experimental studies in spin ice thin films [19, 20], it has been shown that spin ice materials, when put under strain along certain lattice-directions, the system’s physics will change at low temperatures, degeneracies are lifted, and the third law of thermodynamics is restored. Beyond the spin ice models on the pyrochlore lattice, the breathing pyrochlore lattice—a structural variant of the conventional pyrochlore lattice—has emerged as a promising platform for studying novel phases of matter. In this lattice, alternating tetrahedra of different sizes and bond strengths break the inherent symmetry of the system, strongly influencing its magnetic properties [21, 22, 23]. This structural feature has been linked to novel phenomena, including fracton physics and the potential realization of distinct quantum spin liquid phases [24, 25, 26]. This leads to the question of how the anisotropic interactions of a strained system would change the behavior if we allowed for a distinction of the tetrahedral sublattices, as is the case of a breathing pyrochlore lattice. While uniform perturbations, such as external magnetic fields, further neighbor interactions, or global strain, have been extensively explored [27, 17, 18, 28, 29], the effects of spatially dependent modifications remain open.
@@
In this work, we introduce an anisotropic perturbation: a bond-dependent tuning parameter that effectively acts as a sublattice-dependent strain along the $[001]$ direction. This parameter enables independent tuning of the interaction strengths on the two tetrahedral sublattices of the breathing pyrochlore lattice, breaking the lattice symmetry in a controlled and tunable manner. Strain engineering in spin-ice materials, such as Dy2Ti2O7 and Ho2Ti2O7, has been experimentally realized in various studies [20, 19, 30, 17]. However, these studies have primarily focused on uniform strain. Here, we go beyond this by providing a comprehensive theoretical analysis of sublattice-dependent strain on a toy model based on the breathing pyrochlore lattice. Using Monte Carlo simulations, we explore the phases that arise in the ground-state limit for different combinations of strain parameters and analyze the system’s equilibrium properties. Interestingly, we find that in the ground state limit, for different strains, the system effectively decouples into lower-dimensional subsystems, changing the ground state degeneracies. This dimensional reduction can be understood in the framework of intermediate symmetries, where symmetry operations act on lower-dimensional subregions, such as lines or planes, within a higher-dimensional system [31]. While global symmetries act on the whole system and local (gauge) symmetries on single points in space, intermediate symmetries apply uniformly to $d$-dimensional subspaces in a $D$-dimensional system. These symmetries can lead to dimensional reduction through effective energetic decoupling of subsystems [32, 33, 31]. Our work provides a comprehensive understanding of how lattice asymmetry can be leveraged to engineer the ground state and thermodynamic properties of constrained spin systems and generate dimensional reduction.
@@ II Model
We consider a pyrochlore lattice, shown in Fig. 1(a), with interacting spins residing at the lattice sites shown in grey. We assume the system is in the spin ice limit, induced by a large crystal field anisotropy that constrains the spins to point along the local cubic [111] direction [2, 6]. This allows us to simplify the complex quantum mechanical problem of interacting Heisenberg spins on the lattice to a classical problem of antiferromagnetically interacting Ising spins. We additionally assume a distinction between the tetrahedral sublattices, in Fig. 1(a), shown in different colors, which we will refer to as $A$ and $B$ sublattice. This distinction is typically applicable to the breathing pyrochlore lattice [21, 25, 34], where, for example, due to different atoms, the size of the tetrahedra of the two sublattices are different. As a consequence, our interaction constants can be individually tuned on the sublattices. In our model, however, we want to make an additional change to the commonly known spin-ice problem: we will add an additional perturbation that increases or decreases the interaction between two specific spin pairs, $\{S_{0}^{z},S_{3}^{z}\}$ and $\{S_{1}^{z},S_{2}^{z}\}$, by $\delta_{A/B}$ on both sublattices independently and making it thereby anisotropic. This is shown in Fig. 1(a) and (b).
@@
$\begin{split}H&=\sum_{\langle i,j\rangle\in A}J_{A}^{ij}S_{i}^{z}S_{j}^{z}+\sum_{\langle i,j\rangle\in B}J_{B}^{ij}S_{i}^{z}S_{j}^{z},\\ \end{split}$ (1)
with the interaction coefficicent $J_{\alpha}^{ij}=J_{\alpha}+\delta_{\alpha}$ for bonds in the $x,y$-plane, or $ij\in\{03,12\}$, and otherwise just $J_{\alpha}^{ij}=J_{\alpha}$ for all other bonds, as indicated in Fig. 1(b). The chosen numbering of the spins on the tetrahedra is such that the spin pairs with different interaction, $\{S_{0}^{z},S_{3}^{z}\}$ and $\{S_{1}^{z},S_{2}^{z}\}$, are lying in directions $[110]$ and $[-110]$, or, the spins in the $x,y$-plane of the unit cell.
@@
$\begin{split}H=&\frac{J_{A}}{2}\sum_{A}\mathsf{Q}_{A}^{2}+\frac{\delta_{A}}{2}\sum_{A}\left((\mathsf{q}_{A}^{03})^{2}+(\mathsf{q}_{A}^{12})^{2}\right)-2\delta_{A}N_{\text{A-tet}}+\\ &\frac{J_{B}}{2}\sum_{B}\mathsf{Q}_{B}^{2}+\frac{\delta_{B}}{2}\sum_{B}\left((\mathsf{q}_{B}^{03})^{2}+(\mathsf{q}_{B}^{12})^{2}\right)-2\delta_{B}N_{\text{B-tet}}.\end{split}$ (3)
Since we are working with Ising spins, and for computational purposes, we used $|S|=1$. Depending on the length of the spin, the last term of the equation will always read $-\frac{\delta_{\alpha}}{2}N_{\alpha-\text{tet}}4|S|^{2}$, where $N_{\alpha-\text{tet}}=\frac{1}{4}N_{\text{spin}}$ is the number of tetrahedra in the system. We can distinguish 6 different cases by tuning $\delta_{A/B}$ to be either positive, negative, or zero. By diagonalizing the Hamiltonians, we find that the number of ground states on each tetrahedron changes for the different signs of $\delta_{A,B}$. For a single tetrahedron, with $\delta_{\alpha}=\delta=0$, there are $6$ possible ground states on a single tetrahedron; all states that follow the rule “two in - two out.” Introducing the $\delta$-term splits these $6$ states into a group of $4$ and a group of $2$ states. Which of these is the ground state and which is the higher-lying state depends on the sign of $\delta$; see Fig. 1(c). The sign of $\delta$ will also influence the total number of ground states the whole system will have, as well as the possible mobile excitations on top of this ground state. We will introduce the nomenclature listed in Tab. 1 for the six cases inspired by the type of order in the ground state. From now on, we will assume $\delta_{i}<J_{i}$. The phase diagram with the six different phases arising from tuning $\delta_{A}$ and $\delta_{B}$ is shown in Fig. 1(d).
@@ 2. Planar Ice $\delta_{A}>0$ and $\delta_{B}=0$ $(4/3)^{6L_{x}L_{y}L_{z}}$
3. Omni-Plane $\delta_{A}<0$ and $\delta_{B}=0$ $62^{2L_{x}}+2^{2L_{y}}+2^{2L_{z}}-$ 6
@@ 5. Plane-Paramagnetic $\delta_{A}>0$ and $\delta_{B}<0$ $2^{2\cdot L_{z}}$
6. Ferromagnetic $\delta_{A},\delta_{B}<0$ $2$
Table 1: Six limits of the strain parameters with the corresponding ground state degeneracy (GSD). The model is symmetric under the exchange of indices $A$ and $B$. Here, $L_{\alpha}$ is the number of unit cells in one spatial dimension $\alpha\in(x,y,z)$; $N=16\cdot L_{x}\cdot L_{y}\cdot L_{z}$ is the total number of spins in the lattice.
Our goal is to describe the thermodynamic properties of the breathing spin ice in the different strain limits: their theoretical ground state degeneracy as well as experimental signatures, such as the specific heat and entropy density at different temperatures.
@@ III Ground state degeneracy
The number of ground states on a single tetrahedron decreases for any $\delta\neq 0$. In isotropic spin ice, $\delta=0$, we only distinguish between 6 ground states (“two in - two out”), 8 single-monopole states (“three in - one out” or vice versa), and two double-monopoles (“all in” or “all out”). For a single tetrahedron, a total magnetization vector can be defined, $\mathbf{M}_{i}=\sum_{\alpha}S^{z}_{i,\alpha}\mathbf{e}_{\alpha}$, where $\mathbf{e}_{\alpha}$ points along the local $z$ axis of the spin ($\alpha$) position in the tetrahedron ($i$) (see App. B). All six ground states of a single tetrahedron have a non-zero magnetization vector; their magnetization vector will point parallel or antiparallel to one of the three spatial dimensions $(x,y,z)$. Introducing the $\delta$-term energetically splits the six ground states into four lower-lying and two higher-lying states, for $\delta>0$, and two lower-lying and two higher-lying states for $\delta<0$, as indicated in Fig. 1(c).
@@
Figure 2: Tetrahedra-planes and spin-lines. (a) A plane of $A$ tetrahedra (blue) with equal $z$ coordinate. The $A$ tetrahedra and the connecting links on the $B$ tetrahedra are shown in color. The two-dimensional projection onto the plane is shown in black. (b) Two spin lines are shown in color. The 03-line connects spins $S_{0}^{z}$ and $S_{3}^{z}$ in an $x,y$-plane and the 12-line connects spins $S_{1}^{z}$ and $S_{2}^{z}$.
@@
For $\delta>0$, the ground states have in common that the partial charge $\mathsf{q}^{03/12}=0$, so the spin-pairs $S_{0}^{z}=-S_{3}^{z}$ and $S_{1}^{z}=-S_{2}^{z}$ and their magnetization vector points along the $\mathbf{M}_{i}\parallel\hat{x}$ or $\mathbf{M}_{i}\parallel\hat{y}$ direction, while for $\delta<0$, the partial charges $\mathsf{q}^{03}=\mathsf{q}^{12}=\pm 2$ and the opposite, $S_{0}^{z}=S_{3}^{z}$ and $S_{1}^{z}=S_{2}^{z}$, holds, leading to $\mathbf{M}_{i}\parallel\hat{z}$. In any case, however, the total spin sum $\mathsf{Q}_{A/B}=\sum_{i}S_{i}^{z}=0$. From this, we can now build up the ground state for the whole system, matching tetrahedra from the $A$ and the $B$ sublattice. The number of ground states for all systems will decrease compared to spin ice. Importantly, the magnitude of $\delta_{A/B}$ is irrelevant when only considering the ground state degeneracy of the theoretical model. The magnitude will become relevant in experiments as well as when considering thermodynamic properties. Also, the energetic order of possible excitations changes depending on $\delta$, as listed in Tab. 2.
@@
In the following, we will analyze how introducing the $\delta$-terms alters the ground state degeneracy for the whole system. We will first determine how many configurations are ground states on a single tetrahedron of one species ($A/B$) in six different limits and then see how this relates to constructing a ground state on the whole lattice. Importantly, we will also discuss the operator structure necessary to construct all possible ground states. This will then relate to the name chosen for the different phases. The ground state energy for the different limits of the model in Eq. (3) is $E_{\text{gs}}=-\frac{1}{2}(|\delta_{A}|+|\delta_{B}|)N_{\text{spin}}$.
@@
$\mathsf{Q}=0$ $\mathsf{q}^{03}=\mathsf{q}^{12}=0$ $\varepsilon=-2\delta$
$\mathsf{Q}=0$ $\mathsf{q}^{03}=-\mathsf{q}^{12}=\pm 2$ $\varepsilon=2\delta$
$\mathsf{Q}=\pm 2$ $\mathsf{q}^{03}=\pm 2$ or $\mathsf{q}^{12}=\pm 2$ $\varepsilon=2J$
$\mathsf{Q}=\pm 4$ $\mathsf{q}^{03}=\mathsf{q}^{12}=\pm 2$ $\varepsilon=8J+2\delta$

### v2

@@ Abstract
Spin ice systems represent a prime example of constrained spin systems and exhibit rich low-energy physics. In this study, we explore how introducing a tunable anisotropic spin coupling to the conventional Ising spin ice Hamiltonian on the breathing pyrochlore lattice affects the ground state properties of the system. Significant changes are observed in the ground state structure, reflected in the spin structure factor and in a reduction of residual entropy at low temperatures. We theoretically uncover a rich phase diagram by varying the anisotropy and demonstrate how this modification reduces the ground state degeneracy across different phases. Numerical simulations reveal that, at sufficiently low temperatures, the system either undergoes a crossover into a constrained spin ice manifold, characterized by an entropy density that drops below the Pauling entropy of conventional spin ice, or a phase transition into a symmetry-broken state, depending on the perturbations. Additionally, we compute the spin structure factors for the anisotropic model and compare these results to analytical predictions from a self-consistent Gaussian approximation, finding good agreement. This work develops the understanding of spin ice in anisotropic limits, which may be experimentally realized by strain, providing, among others, key signatures in entropy and specific heat.
@@ I Introduction
While the simple nearest-neighbor spin-ice model is by now well understood [2, 6, 7, 8], several questions arise when comparing it to experiments. Experimental studies of Dy2Ti2O7 have revealed, for instance, that thermally equilibrated samples, cooled over long timescales, exhibit an entropy drop below the celebrated Pauling entropy, a phenomenon that remains under active investigation [9, 10]. One possible explanation is the formation of a quantum spin liquid, where the true ground state would be a coherent superposition of the classical configurations, thereby lifting the extensive degeneracy [11, 12, 13]. Another explanation for the release of residual entropy is based on further neighbor or long-range interactions present in spin ice materials. These naturally change the degeneracy of the ground state and lead to an ordered state at low energies [14, 15, 16]. An alternative approach to release the residual entropy has been shown to be through strain engineering. Through theoretical [17, 18] and experimental studies in spin ice thin films [19, 20], it has been shown that spin ice materials, when put under strain along certain lattice directions, the system’s physics will change at low temperatures, degeneracies are lifted, and the third law of thermodynamics is restored. Beyond the spin ice models on the pyrochlore lattice, the breathing pyrochlore lattice—a structural variant of the conventional pyrochlore lattice—has emerged as a promising platform for studying novel phases of matter. In this lattice, alternating tetrahedra of different sizes and bond strengths break the inherent symmetry of the system, strongly influencing its magnetic properties [21, 22, 23]. This structural feature has been linked to novel phenomena, including fracton physics and the potential realization of distinct quantum spin liquid phases [24, 25, 26]. This leads to the question of how the anisotropic interactions of a strained system would change the behavior if we allowed for a distinction of the tetrahedral sublattices, as is the case with a breathing pyrochlore lattice. While uniform perturbations, such as external magnetic fields, further neighbor interactions, or global strain, have been extensively explored [27, 17, 18, 28, 29], the effects of spatially dependent modifications remain open.
@@
In this work, we introduce an anisotropic perturbation: a bond-dependent tuning parameter that effectively acts as a sublattice-dependent strain along the $[001]$ direction. This parameter enables independent tuning of the interaction strengths on the two tetrahedral sublattices of the breathing pyrochlore lattice, breaking the lattice symmetry in a controlled and tunable manner. Strain engineering in spin-ice materials, such as Dy2Ti2O7 and Ho2Ti2O7, has been experimentally realized in various studies [20, 19, 30, 17]. However, these studies have primarily focused on uniform strain. Here, we go beyond this by providing a comprehensive theoretical analysis of sublattice-dependent strain on a toy model based on the breathing pyrochlore lattice. Using Monte Carlo simulations, we explore the phases that arise in the ground-state limit for different combinations of strain parameters and analyze the system’s equilibrium properties. Interestingly, we find that in the ground state limit, for different strains, the system effectively decouples into lower-dimensional subsystems, changing the ground state degeneracies. This dimensional reduction can be understood in the framework of intermediate symmetries, where symmetry operations act on lower-dimensional subregions, such as lines or planes, within a higher-dimensional system [31]. While global symmetries act on the whole system and local (gauge) symmetries on single points in space, intermediate symmetries apply uniformly to $d$-dimensional subspaces in a $D$-dimensional system. These symmetries can lead to dimensional reduction through effective energetic decoupling of subsystems [32, 33, 31].
From a different perspective, this decoupling mechanism is inherently tied to frustration. In the unperturbed spin ice model, all bonds are equally frustrated; introducing anisotropy lifts frustration selectively, favoring certain local configurations and decoupling the system into independent planes or lines. A classical example of this mechanism is found in the triangular lattice Ising antiferromagnet, where increasing anisotropy along one lattice direction reduces frustration and leads to decoupled one-dimensional Ising chains at $T=0$ [34]. By contrast, in unfrustrated three-dimensional systems, anisotropy typically leads only to a crossover in thermodynamic behavior, not full decoupling, since finite coupling persists between subsystems [35, 36, 37]. Thus, frustration plays a crucial role in enabling genuine dimensional reduction in our model. Our work provides a comprehensive understanding of how lattice asymmetry can be leveraged to engineer the ground state and thermodynamic properties of constrained spin systems and generate dimensional reduction in a frustrated system.
@@ II Model
We consider a pyrochlore lattice, shown in Fig. 1(a), with interacting spins residing at the lattice sites shown in grey. We assume the system is in the spin ice limit, induced by a large crystal field anisotropy that constrains the spins to point along the local cubic $[111]$ direction [2, 6]. In rare-earth pyrochlore spin ice materials such as Dy2Ti2O7 and Ho2Ti2O7, the magnetic interactions consist of two main contributions: a short-range superexchange interaction $J_{\text{nn}}$, typically antiferromagnetic, and a long-range dipolar interaction $D_{\text{nn}}$, which is ferromagnetic. Together, these define an effective nearest-neighbor coupling $J_{\text{eff}}=J_{\text{nn}}+D_{\text{nn}}$, which is positive for the materials mentioned and stabilizes the spin ice regime by enforcing the ”2-in, 2-out” ice rule at intermediate temperatures [38, 14]. In this work, we adopt a minimal classical model that retains only the nearest-neighbor exchange interaction, effectively truncating dipolar interactions beyond first neighbors. This simplification maps the system onto a classical Ising problem with antiferromagnetic nearest-neighbor interactions, which remains a good approximation at temperatures above the scale set by the bandwidth introduced by dipolar terms [39, 40]. We additionally assume a distinction between the tetrahedral sublattices, in Fig. 1(a), shown in different colors, which we will refer to as $A$ and $B$ sublattice. This distinction is typically applicable to the breathing pyrochlore lattice [21, 25, 41], where, for example, due to different atoms, the size of the tetrahedra of the two sublattices is different. As a consequence, our interaction constants can be individually tuned on the sublattices. In our model, however, we want to make an additional change to the commonly known spin-ice problem: we will add an additional perturbation that increases or decreases the interaction between two specific spin pairs, $\{S_{0}^{z},S_{3}^{z}\}$ and $\{S_{1}^{z},S_{2}^{z}\}$, by $\delta_{A/B}$ on both sublattices independently and making it thereby anisotropic. This is shown in Fig. 1(a) and (b).
@@
[…]
with the interaction coefficicent $J_{\alpha}^{ij}=J_{\alpha}+\delta_{\alpha}$ for bonds in the $x,y$-plane, or $ij\in\{03,12\}$, and otherwise just $J_{\alpha}^{ij}=J_{\alpha}$ for all other bonds, as indicated in Fig. 1(b). The chosen numbering of the spins on the tetrahedra is such that the spin pairs with different interaction, $\{S_{0}^{z},S_{3}^{z}\}$ and $\{S_{1}^{z},S_{2}^{z}\}$, are lying in directions $[110]$ and [110][-110], or, the spins in the $x,y$-plane of the unit cell.
@@
$\begin{split}H=&\frac{J_{A}}{2}\sum_{A}\mathsf{Q}_{A}^{2}+\frac{\delta_{A}}{2}\sum_{A}\left((\mathsf{q}_{A}^{03})^{2}+(\mathsf{q}_{A}^{12})^{2}\right)+\\ &\frac{J_{B}}{2}\sum_{B}\mathsf{Q}_{B}^{2}+\frac{\delta_{B}}{2}\sum_{B}\left((\mathsf{q}_{B}^{03})^{2}+(\mathsf{q}_{B}^{12})^{2}\right)\\ &-2(J_{A}+\delta_{A})N_{\text{A-tet}}-2(J_{B}+\delta_{B})N_{\text{B-tet}}.\end{split}$ (3)
Since we are working with Ising spins, and for computational purposes, we used $|S|=1$. Depending on the length of the spin, the last term of the equation will always read $\frac{(J_{\alpha}+\delta_{\alpha})}{2}N_{\alpha-\text{tet}}4|S|^{2}$, where $N_{\alpha-\text{tet}}=\frac{1}{4}N_{\text{spin}}$ is the number of tetrahedra in the system. From here on, we will neglect this constant energy shift and set the last term in eq. (3) to zero. We can distinguish six different cases by tuning $\delta_{A/B}$ to be either positive, negative, or zero. The convenient form of the Hamiltonian in Eq. (3) makes the ground-state structure apparent: since we assume $J_{A/B}>0$, a ground state must fulfill $\mathsf{Q}_{A/B}=0$. However, depending on the sign of $\delta_{A/B}$ the energy is minimized by either minimizing $\mathsf{q}_{A/B}^{2}$ – corresponding to $\mathsf{q}_{A/B}=0$ – or maximizing the latter by $\mathsf{q}_{A/B}=\pm 2$.
For a single tetrahedron, with $\delta_{\alpha}=\delta=0$, there are $6$ possible ground states on a single tetrahedron; all states that follow the rule “two in - two out” and have $\mathsf{Q}_{A/B}=0$. Introducing the $\delta$-term splits these $6$ states into a group of $4$ ($\mathsf{q}_{A/B}=0$ ) and a group of $2$ ($\mathsf{q}_{A/B}=\pm 2$) states. Which of these is the ground state and which is the higher-lying state depends on the sign of $\delta$; see Fig. 1(c).
The sign of $\delta$ will also influence the total number of ground states the whole system will have, as well as the possible mobile excitations on top of this ground state. We will introduce the nomenclature listed in Tab. 1 for the six cases inspired by the type of order in the ground state. From now on, we will assume $\delta_{i}<J_{i}$. The phase diagram with the six different phases arising from tuning $\delta_{A}$ and $\delta_{B}$ is shown in Fig. 1(d).
@@ 2. Planar Ice $\delta_{A}>0$ and $\delta_{B}=0$ $(4/3)^{6L_{x}L_{y}L_{z}}$
3. Omni-Plane $\delta_{A}<0$ and $\delta_{B}=0$ $2^{2L_{x}}+2^{2L_{y}}+2^{2L_{z}}-6$
@@ 5. Plane-Paramagnetic $\delta_{A}>0$ and $\delta_{B}<0$ $2^{2\cdot L_{z}}$
6. Ferromagnetic $\delta_{A},\delta_{B}<0$ 22
Table 1: Six limits of the strain parameters with the corresponding ground state degeneracy (GSD). The model is symmetric under the exchange of indices AA and $B$. Here, $L_{\alpha}$ is the number of unit cells in one spatial dimension $\alpha\in(x,y,z)$; $N=16\cdot L_{x}\cdot L_{y}\cdot L_{z}$ is the total number of spins in the lattice.
Our goal is to describe the thermodynamic properties of the breathing spin ice in the different strain limits: their theoretical ground state degeneracy and mobility of excitations on top of the ground state, as well as experimental signatures, such as the specific heat and entropy density at different temperatures.
@@ III Ground state degeneracy
The number of ground states on a single tetrahedron decreases for any $\delta\neq 0$. In isotropic spin ice, $\delta=0$, we only distinguish between 6 ground states (“two in - two out”), 8 single-monopole states (“three in - one out” or vice versa), and two double-monopoles (“all in” or “all out”). For a single tetrahedron, a total magnetization vector can be defined, $\mathbf{M}_{i}=\sum_{\alpha}S^{z}_{i,\alpha}\mathbf{e}_{\alpha}$, where $\mathbf{e}_{\alpha}$ points along the local $z$ axis of the spin ($\alpha$) position in the tetrahedron ($i$) (see App. C). All six ground states of a single tetrahedron have a non-zero magnetization vector; their magnetization vector will point parallel or antiparallel to one of the three spatial dimensions $(x,y,z)$. Introducing the $\delta$-term energetically splits the six ground states into four lower-lying and two higher-lying states, for $\delta>0$, and two lower-lying and two higher-lying states for $\delta<0$, as indicated in Fig. 1(c).
@@
Figure 2: Tetrahedra-planes and spin-lines. (a) A plane of tetrahedra of the same kind is shown. Here, AA tetrahedra (blue) with equal $z$ coordinate, or an $x,y$-plane, are depicted as an example. The $A$ tetrahedra and the connecting links on the $B$ tetrahedra (lavender) are shown in color. The two-dimensional projection onto the plane is shown in black. Planes can be formed perpendicular to all spatial directions by either $A$ or $B$ tetrahedra. (b) Two spin lines are shown in color. The 03-line connects spins $S_{0}^{z}$ and $S_{3}^{z}$ in an $x,y$-plane and the 12-line connects spins $S_{1}^{z}$ and $S_{2}^{z}$.
@@
For $\delta>0$, the ground states have in common that the partial charge $\mathsf{q}^{03/12}=0$, so the spin-pairs $S_{0}^{z}=-S_{3}^{z}$ and $S_{1}^{z}=-S_{2}^{z}$ and their magnetization vector points along the $\mathbf{M}_{i}\parallel\hat{x}$ or $\mathbf{M}_{i}\parallel\hat{y}$ direction, while for $\delta<0$, the partial charges $\mathsf{q}^{03}=\mathsf{q}^{12}=\pm 2$ and the opposite, $S_{0}^{z}=S_{3}^{z}$ and $S_{1}^{z}=S_{2}^{z}$, holds, leading to $\mathbf{M}_{i}\parallel\hat{z}$. In any case, however, the total spin sum $\mathsf{Q}_{A/B}=\sum_{i}S_{i}^{z}=0$. From this, we can now build up the ground state for the whole system, matching tetrahedra from the $A$ and the $B$ sublattices. The number of ground states for all systems will decrease compared to spin ice. Importantly, the magnitude of $\delta_{A/B}$ is irrelevant when only considering the ground state degeneracy of the theoretical model. The magnitude will become relevant in experiments as well as when considering thermodynamic properties. Also, the energetic order of possible excitations changes depending on $\delta$, as listed in Tab. 2.
@@
In the following, we will analyze how introducing the $\delta$-terms alters the ground state degeneracy for the whole system. We will first determine how many configurations are ground states on a single tetrahedron of one species ($A/B$) in six different limits and then see how this relates to constructing a ground state on the whole lattice. Importantly, we will also discuss the operator structure necessary to construct all possible ground states. This will then relate to the name chosen for the different phases. Lastly, we will also mention the mobility of excitations on top of the ground state. This will make apparent a connection to fracton physics, characterized by quasiparticle excitations being restricted in their mobility [44, 45, 46, 47, 48, 24, 25].
@@
$\mathsf{Q}=0$ $\mathsf{q}^{03}=\mathsf{q}^{12}=0$ $\varepsilon=0$
$\mathsf{Q}=0$ $\mathsf{q}^{03}=-\mathsf{q}^{12}=\pm 2$ $\varepsilon=4\delta$
$\mathsf{Q}=\pm 2$ $\mathsf{q}^{03}=\pm 2$ or $\mathsf{q}^{12}=\pm 2$ $\varepsilon=2J+2\delta$
$\mathsf{Q}=\pm 4$ $\mathsf{q}^{03}=\mathsf{q}^{12}=\pm 2$ $\varepsilon=8J+4\delta$